# The viscoelastic rheology of transient diffusion creep

### v1

@@ H.4 Andrade
The Andrade model has compliance and creep function
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Gamma(1+\alpha)}{(s\tau_{A})^{\alpha}}+\frac{1}{s\tau_{M}}}\right),$ (47)
$J(t)$ $=J_{0}\left(1+\left(\frac{t}{\tau_{A}}\right)^{\alpha}+\frac{t}{\tau_{M}}\right),$ (48)
@@ H.5 Extended Burgers
$s\tilde{J}(s)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\frac{\gamma\rho(\gamma)}{s+\gamma}\;\mathrm{d}\gamma+\frac{1}{s\tau_{M}}}\right),$ (49)
$J(t)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\left(1-\mathrm{e}^{-\gamma t}\right)\rho(\gamma)\;\mathrm{d}\gamma+\frac{t}{\tau_{M}}}\right),$ (50)
@@
$\int_{0}^{\infty}\rho(\gamma)\;\mathrm{d}\gamma=1.$ (51)
The Burgers model in (45) is a special case of (49) with a single frequency, $\rho(\gamma)=\delta(\gamma-1/\tau_{r})$, where $\delta(x)$ is the Dirac delta function. The integral in (49) can be recognised as the Stieltjes transform of $\gamma\rho(\gamma)$.
The extended Burgers model in (49) is sufficiently general to describe any linear viscoelastic response. A commonly-used parametrisation of $\rho(\gamma)$ has a power-law distribution in a period band between a short-period cut-off $\tau_{L}$ and a long-period cut-off $\tau_{H}$ (e.g. [41, 19]). In the present work we will neglect the short-period cut-off, and write the density as
$\rho(\gamma)=\begin{cases}\dfrac{\alpha\tau_{H}}{(\gamma\tau_{H})^{\alpha+1}},\quad\quad&\gamma\tau_{H}>1,\\ 0,\quad\quad&\text{otherwise.}\end{cases}$ (52)
in terms of the power law exponent $\alpha$. The Stieltjes transform can be determined analytically, and (49) can be written as
$s\tilde{J}(s)=J_{0}\left({1+\Delta\,{}_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)+\frac{1}{s\tau_{M}}}\right)$ (53)
@@
$J(t)=J_{0}\left(1+\Delta\left(1-\alpha\,E_{1+\alpha}\left(\frac{t}{\tau_{H}}\right)\right)+\frac{t}{\tau_{M}}\right)$ (54)
@@
${}_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)\sim\frac{\Gamma(1+\alpha)\Gamma(1-\alpha)}{(s\tau_{H})^{\alpha}}$ (55)
@@
$\tau_{A}=\tau_{H}\left(\Gamma(1-\alpha)\,\Delta\right)^{-1/\alpha}.$ (56)
@@
$s\tilde{J}(s)\sim J_{0}\left({1+\Delta+\frac{1}{s\tau_{M}}}\right)$ (57)
@@ Appendix I Voigt averaging
$s\tilde{G}^{V}(s)=\frac{2}{5}s\tilde{G}^{(1)}(s)+\frac{3}{5}s\tilde{G}^{(2)}(s),$ (58)
@@
$s\tilde{G}(s)\sim s\eta-\frac{s^{2}\eta^{2}}{G_{ss}}+\cdots$ (59)
@@
$\frac{(\eta^{V})^{2}}{G_{ss}^{V}}=\frac{2}{5}\frac{(\eta^{(1)})^{2}}{G_{ss}^{(1)}}+\frac{3}{5}\frac{(\eta^{(2)})^{2}}{G_{ss}^{(2)}}.$ (61)
@@
$s\tilde{G}(s)\sim G_{0}-\frac{G_{0}\Gamma(1+\alpha)}{\left(s\tau_{A}\right)^{\alpha}}+\cdots$ (62)
@@
$\frac{G_{0}^{V}}{\left(\tau_{A}^{V}\right)^{\alpha}}=\frac{2}{5}\frac{G_{0}^{(1)}}{\left(\tau_{A}^{(1)}\right)^{\alpha}}+\frac{3}{5}\frac{G_{0}^{(2)}}{\left(\tau_{A}^{(2)}\right)^{\alpha}}.$ (64)
@@
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}\frac{w_{j}f_{j}}{\nu-\nu_{j}}}{\displaystyle\sum_{j=1}^{m}\frac{w_{j}}{\nu-\nu_{j}}}$ (65)
@@
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}w_{j}f_{j}\prod_{k\neq j}(\nu-\nu_{k})}{\displaystyle\sum_{j=1}^{m}w_{j}\prod_{k\neq j}(\nu-\nu_{k})},$ (66)
from which it is clear the function is the ratio of two polynomials of degree $m-1$. The weights can all be scaled by a constant factor and the rational function will be unchanged. To specify the weights uniquely, here we fix $w_{1}=1$.
@@
Table 2 gives parameters which fit the data for hexagons shown in Figure 1. The fits have just three free parameters, and represent the properties as a ratio of two linear functions of $\nu$ (i.e. $m=2$). The fits for the scaled moduli $G_{0}/\mu$ and $G_{ss}/\mu$ have relative errors $<10^{-6}$ which suggests that the representation of these quantities in terms of a ratio of linear functions may in fact be exact, similar to the analytic result for $G_{0}/\mu$ for spheres by Zener [42]. Larger errors are seen in fitting the scaled Andrade time $\tau_{A}/\tau_{M}$, but the parameters in the table still represent an excellent fit with relative errors $<10^{-3}$.
@@
Table 3 gives parameters for fits of the data plotted in Figure 4 for tetrakaidecahedral grains. Unlike the case of hexagons, the data require higher order polynomials to be fit well; the table uses a fit of a ratio of quadratics ($m=3$), which fits the scaled moduli to within a relative error of $<10^{-3}$, and the scaled Andrade times $\tau_{A}/\tau_{M}$ to within $<2\times 10^{-2}$.
@@
The scaling argument given here follows that in Lee et al. [6]. For material deforming at a typical strain $\epsilon$, the energy stored elastically per unit volume scales as
$\mathcal{U}\sim\mu\epsilon^{2}.$ (67)
@@
$\mathcal{D}\sim\frac{\Omega\delta D^{\text{gb}}}{kT\omega V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S=\frac{l^{3}}{\mu V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S$ (68)
where $l$ is the lengthscale defined in (5).
With a stress singularity at the triple junction, the normal stress scales as
$p\sim\mu\epsilon\left(r/d\right)^{-\lambda}$ (69)
@@
$\nabla_{\perp}p\sim\frac{\mu\epsilon}{d}\left(r/d\right)^{-(\lambda+1)}.$ (70)
The surface integral term in (68) can be approximated in 2D as
$\int_{l}^{\infty}|\nabla_{\perp}p|^{2}\,\mathrm{d}r\sim\frac{\mu^{2}\epsilon^{2}}{d}\left(l/d\right)^{-(2\lambda+1)}$ (71)
where it is assumed that diffusion smooths the singularity within a distance $l$ of the triple junction, and that $l\ll d$. In 2D the grain volume $V\sim d^{2}$ and hence (68) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-2(1-\lambda)/3}.$ (72)
@@
$Q^{-1}\sim\frac{\mathcal{D}}{\mathcal{U}}\sim\left(\omega\tau_{M}\right)^{-\alpha}$ (73)
@@
$\alpha=\frac{2}{3}\left(1-\lambda\right).$ (74)
@@
The main focus of this work has been on the process of grain boundary diffusion, as that process has been suggested to be the one of relevance in a number of experimental studies [e.g. 2]. However, there are circumstances where diffusion within the interiors of grains dominates over diffusion on the grain boundaries, and in steady-state this is known as Nabarro-Herring creep. From a modelling perspective volume diffusion can be treated in a very similar way to surface diffusion, and simply results in replacing a surface integral in the weak form of the equations by a volume integral, i.e. instead of (33), one has
$-\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega D}{kTs^{*}}\int_{V_{\text{grain}}}\nabla p^{*}\cdot\nabla q\;\mathrm{d}V=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (75)
@@
$\eta=\frac{5}{288}\frac{kTd^{2}}{D\Omega},$ (76)
@@
$l=\left(\frac{\mu D\Omega}{kT\omega}\right)^{1/2},$ (77)
which replaces the surface diffusion result in (5). The expression for dissipation in (68) becomes
$\mathcal{D}\sim\frac{\Omega D}{kT\omega V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V=\frac{l^{2}}{\mu V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V$ (78)
and the expression in (72) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-(1-\lambda)},$ (79)
and hence the expression relating the Andrade exponent to the stress exponent in (74) becomes
$\alpha=1-\lambda.$ (80)
@@ Acknowledgments
I thank H. Yamauchi and Y. Takei for kindly sharing their data. I also thank H. Innes, G. Johnson, T. Breithaupt, D. Al-Attar, Y. Takei and D. McKenzie for insightful discussions on transient creep.

### v2

@@ H.4 Andrade
The Andrade model [44] has compliance and creep function
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Gamma(1+\alpha)}{(s\tau_{A})^{\alpha}}+\frac{1}{s\tau_{M}}}\right),$ (H 47)
$J(t)$ $=J_{0}\left(1+\left(\frac{t}{\tau_{A}}\right)^{\alpha}+\frac{t}{\tau_{M}}\right),$ (H 48)
@@ H.5 Extended Burgers
$s\tilde{J}(s)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\frac{\gamma\rho(\gamma)}{s+\gamma}\;\mathrm{d}\gamma+\frac{1}{s\tau_{M}}}\right),$ (H 49)
$J(t)$ $=J_{0}\left({1+\Delta\int_{0}^{\infty}\left(1-\mathrm{e}^{-\gamma t}\right)\rho(\gamma)\;\mathrm{d}\gamma+\frac{t}{\tau_{M}}}\right),$ (H 50)
@@
$\int_{0}^{\infty}\rho(\gamma)\;\mathrm{d}\gamma=1.$ (H 51)
The Burgers model in (H 45) is a special case of (H 49) with a single frequency, $\rho(\gamma)=\delta(\gamma-1/\tau_{r})$, where $\delta(x)$ is the Dirac delta function. The integral in (H 49) can be recognised as the Stieltjes transform of $\gamma\rho(\gamma)$.
The extended Burgers model in (H 49) is sufficiently general to describe any linear viscoelastic response. A commonly-used parametrisation of $\rho(\gamma)$ has a power-law distribution in a period band between a short-period cut-off $\tau_{L}$ and a long-period cut-off $\tau_{H}$ (e.g. [45, 19]). In the present work we will neglect the short-period cut-off, and write the density as
$\rho(\gamma)=\begin{cases}\dfrac{\alpha\tau_{H}}{(\gamma\tau_{H})^{\alpha+1}},\quad\quad&\gamma\tau_{H}>1,\\ 0,\quad\quad&\text{otherwise.}\end{cases}$ (H 52)
in terms of the power law exponent $\alpha$. The Stieltjes transform can be determined analytically, and (H 49) can be written as
$s\tilde{J}(s)=J_{0}\left({1+\Delta\,{}_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)+\frac{1}{s\tau_{M}}}\right)$ (H 53)
@@
$J(t)=J_{0}\left(1+\Delta\left(1-\alpha\,E_{1+\alpha}\left(\frac{t}{\tau_{H}}\right)\right)+\frac{t}{\tau_{M}}\right)$ (H 54)
@@
$_{2}F_{1}\left(1,\alpha,1+\alpha,-s\tau_{H}\right)\sim\frac{\Gamma(1+\alpha)\Gamma(1-\alpha)}{(s\tau_{H})^{\alpha}}$ (H 55)
@@
$\tau_{A}=\tau_{H}\left(\Gamma(1-\alpha)\,\Delta\right)^{-1/\alpha}.$ (H 56)
@@
$s\tilde{J}(s)\sim J_{0}\left({1+\Delta+\frac{1}{s\tau_{M}}}\right)$ (H 57)
@@ Appendix I Voigt averaging
$s\tilde{G}^{V}(s)=\frac{2}{5}s\tilde{G}^{(1)}(s)+\frac{3}{5}s\tilde{G}^{(2)}(s),$ (I 58)
@@
$s\tilde{G}(s)\sim s\eta-\frac{s^{2}\eta^{2}}{G_{ss}}+\cdots$ (I 59)
@@
$\frac{(\eta^{V})^{2}}{G_{ss}^{V}}=\frac{2}{5}\frac{(\eta^{(1)})^{2}}{G_{ss}^{(1)}}+\frac{3}{5}\frac{(\eta^{(2)})^{2}}{G_{ss}^{(2)}}.$ (I 61)
@@
$s\tilde{G}(s)\sim G_{0}-\frac{G_{0}\Gamma(1+\alpha)}{\left(s\tau_{A}\right)^{\alpha}}+\cdots$ (I 62)
@@
$\frac{G_{0}^{V}}{\left(\tau_{A}^{V}\right)^{\alpha}}=\frac{2}{5}\frac{G_{0}^{(1)}}{\left(\tau_{A}^{(1)}\right)^{\alpha}}+\frac{3}{5}\frac{G_{0}^{(2)}}{\left(\tau_{A}^{(2)}\right)^{\alpha}}.$ (I 64)
@@
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}\frac{w_{j}f_{j}}{\nu-\nu_{j}}}{\displaystyle\sum_{j=1}^{m}\frac{w_{j}}{\nu-\nu_{j}}}$ (J 65)
@@
$r(\nu)=\frac{\displaystyle\sum_{j=1}^{m}w_{j}f_{j}\prod_{k\neq j}(\nu-\nu_{k})}{\displaystyle\sum_{j=1}^{m}w_{j}\prod_{k\neq j}(\nu-\nu_{k})},$ (J 66)
from which it is clear the function is the ratio of two polynomials of degree m1m-1. The weights can all be scaled by a constant factor and the rational function will be unchanged. To specify the weights uniquely, here we fix $w_{1}=1$.
@@
Table 2 gives parameters which fit the data for hexagons shown in Figure 1. The fits have just three free parameters, and represent the properties as a ratio of two linear functions of $\nu$ (i.e. $m=2$). The fits for the scaled moduli $G_{0}/\mu$ and $G_{ss}/\mu$ have relative errors $10^{-6}$ which suggests that the representation of these quantities in terms of a ratio of linear functions may in fact be exact, similar to the analytic result for $G_{0}/\mu$ for spheres by Zener [46]. Larger errors are seen in fitting the scaled Andrade time $\tau_{A}/\tau_{M}$, but the parameters in the table still represent an excellent fit with relative errors $10^{-3}$.
@@
Table 3 gives parameters for fits of the data plotted in Figure 4 for tetrakaidecahedral grains. Unlike the case of hexagons, the data require higher order polynomials to be fit well; the table uses a fit of a ratio of quadratics ($m=3$), which fits the scaled moduli to within a relative error of $10^{-3}$, and the scaled Andrade times $\tau_{A}/\tau_{M}$ to within $2\times 10^{-2}$.
@@
The scaling argument given here follows that in Lee et al. [7]. The balance of terms in (A 17) determines a natural lengthscale $l$ over which grain boundary diffusion acts when forced at a frequency $\omega$. For a typical strain $\epsilon$, the surface tractions are of magnitude $p\sim\mu\epsilon$, and jumps in the normal velocity of magnitude $\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]\sim\epsilon l\omega$. Substituting these scalings into (A 17), with the surface Laplacian scaling as $1/l^{2}$, yields the expression for $l$ given in (5).
The energy stored elastically per unit volume scales as
$\mathcal{U}\sim\mu\epsilon^{2}.$ (K 67)
@@
$\mathcal{D}\sim\frac{\Omega\delta D^{\text{gb}}}{kT\omega V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S=\frac{l^{3}}{\mu V}\int_{S_{i}}|\nabla_{\perp}p|^{2}\;\mathrm{d}S.$ (K 68)
With a stress singularity at the triple junction, the normal stress is
$p\sim\mu\epsilon\left(r/d\right)^{-\lambda}$ (K 69)
@@
$\nabla_{\perp}p\sim\frac{\mu\epsilon}{d}\left(r/d\right)^{-(\lambda+1)}.$ (K 70)
The surface integral term in (K 68) can be approximated in 2D as
$\int_{l}^{\infty}|\nabla_{\perp}p|^{2}\,\mathrm{d}r\sim\frac{\mu^{2}\epsilon^{2}}{d}\left(l/d\right)^{-(2\lambda+1)}$ (K 71)
where it is assumed that diffusion smooths the singularity within a distance $l$ of the triple junction, and that $l\ll d$. In 2D the grain volume $V\sim d^{2}$ and hence (K 68) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-2(1-\lambda)/3}.$ (K 72)
@@
$Q^{-1}\sim\frac{\mathcal{D}}{\mathcal{U}}\sim\left(\omega\tau_{M}\right)^{-\alpha}$ (K 73)
@@
$\alpha=\frac{2}{3}\left(1-\lambda\right).$ (K 74)
@@
The main focus of this work has been on the process of grain boundary diffusion, as that process has been suggested to be the one of relevance in a number of experimental studies [e.g. 2]. However, there are circumstances where diffusion within the interiors of grains dominates over diffusion on the grain boundaries, and in steady-state this is known as Nabarro-Herring creep. From a modelling perspective volume diffusion can be treated in a very similar way to surface diffusion, and simply results in replacing a surface integral in the weak form of the equations by a volume integral, i.e. instead of (F 33), one has
$\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega D}{kTs^{*}}\int_{V_{\text{grain}}}\nabla p^{*}\cdot\nabla q\;\mathrm{d}V=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (L 75)
@@
$\eta=\frac{5}{288}\frac{kTd^{2}}{D\Omega},$ (L 76)
@@
$l=\left(\frac{\mu D\Omega}{kT\omega}\right)^{1/2},$ (L 77)
which replaces the surface diffusion result in (5). The expression for dissipation in (K 68) becomes
$\mathcal{D}\sim\frac{\Omega D}{kT\omega V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V=\frac{l^{2}}{\mu V}\int_{V}|\nabla p|^{2}\;\mathrm{d}V$ (L 78)
and the expression in (K 72) becomes
$\mathcal{D}\sim\mu\epsilon^{2}\left(l/d\right)^{2(1-\lambda)}\sim\mu\epsilon^{2}\left(\omega\tau_{M}\right)^{-(1-\lambda)},$ (L 79)
and hence the expression relating the Andrade exponent to the stress exponent in (K 74) becomes
$\alpha=1-\lambda.$ (L 80)
@@ Acknowledgments
I thank H. Yamauchi and Y. Takei for kindly sharing their data. I thank two anonymous reviewers for their helpful comments. I also thank H. Innes, G. Johnson, T. Breithaupt, D. Al-Attar, Y. Takei and D. McKenzie for insightful discussions on transient creep.